### v1

@@ 7 Discussion
Hence, the Criss-Cross rule is fixed $(r,s)$ in $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$, and $(s,r)$ is a valid reverse Criss-Cross rule pivot.
Case 2. Let the part (b) hold. Similar to the first case, applying the Criss-Cross rule on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ the first step fix at $s\in\tilde{N}_{\neq g}$ and second step pinpoint $r$ in the column of $\mathbf{x}_{s}$, it follows $(s,r)$ is still valid reverse Criss-Cross rule pivot.
@@
Table 8 shows that the complemented Criss-Cross rule can find all vertices using the same hypercubes in Table 2.
$\mathbb{R}^{4}$
$\mathbb{R}^{5}$
$\mathbb{R}^{6}$
$\mathbb{R}^{7}$
$\mathbb{R}^{8}$
$\mathbb{R}^{9}$
$\mathbb{R}^{10}$
Truth
16
32
64
128
256
512
1024
After Complement
16
32
64
128
256
512
1024
Original
5
6
7
8
9
10
11
Table 8: The result from complete and incomplete reverse Criss-Cross rule.
Compared with the if condition in [4], our complement exerts two more restrictions. But the complexity of examining the valid reverse of the Criss-Cross rule is still $\mathcal{O}(n)$. Therefore, the total complexity is $\mathcal{O}(n^{2}dv)$. In contrast, the complexity of the Zero rule can be as low as $\mathcal{O}(nd^{4}v)$, which means that the Zero rule is suitable for an arrangement with more hyperplanes, while the Criss-Cross rule is appropriate for an arrangement with higher dimensions.
8 Conclusion
In this paper, we present the Zero rule for the VE problem of hyperplane arrangements. Compared to the classical Criss-Cross rule, the Zero rule enjoys several computationally friendly properties: i) It can get rid of objective functions, which saves the computation. ii) The condition between the Zero rule and its valid reverse is sufficient and necessary, which prevents the algorithm from going into a loop and catastrophically skipping legitimate dictionaries. iii) In $\mathbb{R}^{d}$, its pivot steps are at most $d$, which can transform the unpredictable depth into a predictable maximum depth of $d$. iv) Its termination dictionary within the entire arrangement is unique. This approach, originating from the arrangement of hyperplanes itself, also contributes to establishing rules with fewer convergence points on a polytope. Because of these properties, the complexity of the VE algorithm using the Zero can be as low as $\mathcal{O}(nd^{4}v)$. Systematic and comprehensive experiments have confirmed that the Zero rule is the state of the art in the VE problem of hyperplane arrangements. In the future, more efforts will be made to increase the proportion of optimal solutions in the dictionary or reduce the amount of information needed to verify reversibility to further escalate the algorithmic efficiency.
9 Acknowledgement

### v2

@@ 8 Acknowledgement
Hence, the Criss-Cross rule is fixed $(r,s)$ in $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$, and $(s,r)$ is a valid reverse Criss-Cross pivot.
Case 2. Let the part (b) hold. Similar to the first case, applying the Criss-Cross rule on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ the first step fix at $s\in\tilde{N}_{\neq g}$ and second step pinpoint $r$ in the column of $\mathbf{x}_{s}$, it follows $(s,r)$ is still valid reverse Criss-Cross pivot.
@@
Figure 13: Illustration of the Proposition 8 to the valid reverse Criss-Cross pivot.
By applying this condition, checking whether a pivot is a valid reverse Criss-Cross pivot can be simplified as follows: if the necessary and sufficient conditions are met, then it is a valid reverse Criss-Cross pivot. Thus update the dictionary and reset the check position. Conversely, if the conditions are not met, one can proceed to check the next position in the current dictionary. The modification is similar with the RHS of Figure 5. The only difference is that the rule checked here is the Criss-Cross rule rather than the Zero rule. A graphic explanation of the Enhanced AF is shown in the LHS of Figure 6, and the pseudocode under this formulation is shown in Algorithm 2.
Based on the necessary and sufficient conditions and Figure 13, it can be observed that the necessary and sufficient conditions require checking at most $2(2d+2(n-d))=4n$ positions. Thus the analysis for Enhanced AF is relatively straightforward. Regardless of whether the pivots are valid or not, it consumes at most $4n\in\mathcal{O}(n)$ operations at each position in the process similar with the RHS of Figure 5, the total operations that the algorithm Enhanced AF requires for testing valid reverse pivots are $\mathcal{O}(n^{2}dv)$. Thus, the total operations cost in checking and performing valid reverse pivot are:
$\mathcal{O}(n^{2}dv+ndv)=\mathcal{O}(n^{2}dv)$ (17)
As a result, comparing Eq. (8) with Eq. (17), it can be seen that the total cost of the reverse search on each dictionary by the Enhanced AF is always better than the original AF.
Algorithm 2 Enhanced AF, reverse search of an optimal dictionary
function $\mathtt{Search}(B,N,\bar{A})$
     $i=1,j=2$;Use $i,j$ represent the $i$-th row, $j$-th column in $\bar{A}$.
     if $\mathtt{lex-min}(B,N,\bar{A})$==1 then
         Print $B$;
     end if
     while $j\leq\mathtt{length}(N)+1$ do
         if $j\leq\mathtt{length}(N)$ then
              if $\mathtt{reverse}(B,N,\bar{A},i,j)==1$ then % A valid reverse pivot is found.
                  $[B,N,\bar{A}]=\mathtt{pivot}(B,N,\bar{A},i,j)$;Compute and update dictionary.
                  if $\mathtt{lex-min}(B,N,\bar{A})==1$ then
                       Print $B$;
                  end if
                  $i=1,j=2;$Initialize the position.
              else
                  $[i,j]=\mathtt{increment}(i,j)$;Invalid, go next position.
              end if
         elseEach position on this dictionary has been checked.
              $[i,j]=\mathtt{select}(B,N,\bar{A})$;Use Criss-Cross rule to select a position.
              if both $i$ and $j$ nonempty thenThe Criss-Cross rule find a position.
                  $[B,N,\bar{A}]=\mathtt{pivot}(B,N,\bar{A},i,j)$;Back to parent dictionary.
                  $[i,j]=\mathtt{find}(i,j)$;Find the corresponding position.
                  $[i,j]=\mathtt{increment}(i,j)$;
              elseThe Criss-Cross rule does not find a position.
                  Break;The function $Search$ terminated.
              end if
         end if
     end while
end function function $\mathtt{lex-min}(B,N,\bar{A})$Test if the current basis output.
function $\mathtt{reverse}(B,N,\bar{A},i,j)$Test if pivot$(B(i),N(j))$ a valid reverse pivot.
function $[B,N,\bar{A}]=\mathtt{pivot}(B,N,\bar{A},i,j)$Compute the new dictionary.
function $[i,j]=\mathtt{increment}(i,j)$Go next position on the dictionary.
function $[i,j]=\mathtt{select}(i,j)$Use the Criss-Cross rule to find a position.
function $[i,j]=\mathtt{find}(i,j)$Find corresponding position after Criss-Cross pivot.